# MonoBox: Tightness-free Box-supervised Polyp Segmentation using Monotonicity Constraint

### v4

@@ 4.3 Comparison with Anti-noise SOTAs
We view MonoBox as a plug-and-play anti-noise method, and compare it with four state-of-the-art (SOTA) anti-noise methods [40, 21, 34, 44] on the Real noisy dataset and the Synthetic noisy dataset with noise level $\sigma=0.2$. We train these methods using two box-supervised segmentation backbones, i.e., WeakPolyp [33] and IBoxCLA [36]. Note that, we denote UB and LB as backbones trained using the clean and noisy datasets, respectively, and since there are no clean annotations in the Real noisy dataset, we do not report results of UB for the Real noisy dataset.
The comparison results are provided in Table 1. Comparing the results of LB and UB of the two backbones, we can see that on Synthetic noisy dataset, the noise leads a decrease of Dice by 7.4% and 9.2% for WeakPolyp and IBoxCLA, respectively. This exhibits the high sensitivity of the existing box-supervised segmentation methods to noisy boxes. When using MonoBox for optimization, the performance margin compared to the UB is significantly narrowed for both backbones. Specifically, Dice of IBoxCLA is improved by 4.6% and 6.8% , and Dice of WeakPolyp is improved by 2.9% and 6.3%, on the Real and Synthetic noisy datasets, respectively.
Moreover, our method achieves the best performance among all anti-noise methods. IBoxCLA with our method exceeds that with the second-best method, i.e., FSRM [44], by 3.3% and 5.5% in Dice on the Real and Synthetic noisy datasets, respectively. WeakPolyp with our method outperforms that with FSRM by 1.9% and 4.1% in Dice on the Real and Synthetic noisy datasets, respectively. This is because that FSRM aims to correct the original noisy boxes using the model’s predictions, but the predictions guided by traditional MIL are unreliable under the supervision of noisy labels. In contrast, the monotonicity constraint of MonoBox can provide reliable supervision even under noisy labels.
Fig. 3 shows the visualization results of different methods using IBoxCLA as the backbone. It can be seen that the model trained with our method can accurately find the inconspicuous polyps and segment more reasonable boundaries.
4.4 Ablation Study
4.4.1 Effectiveness of the key components.
@@ 4.4 Ablation Study
To verify the effectiveness two key components of MonoBox, i.e., monotonicity constraint (MC) and label correction (LC), we train three variants of MonoBox by disabling MC and/or LC, and IBoxCLA is used as the box-supervised segmentation backbone. The segmentation results are presented in Table 2 and Fig. 4. Based on these results, three key conclusions can be made as follows:
@@
(1) By comparing the first two rows in Table 2 corresponding to (a) and (b) in Fig. 4. MC improves the tolerance to noisy box annotations, leading to an increase of 3.5%, 4.2% and a decrease of 0.359 px on the Real noisy dataset, and a increase of 5.2%, 7.2% and a decrease of 0.655 px on the Synthetic noisy dataset in terms of Dice, IoU, and HD, respectively. From Fig. 4(a), we can see that on the Real noisy dataset, the baseline tends to produce small thorn-like regions (indicated by yellow arrows) at the boundary. This is because the doctors often annotated over-sized boxes, and thus the model has to output splinters to touch the incorrect boundaries. This phenomenon does not occur in the Synthetic noisy dataset, because we use gaussian distributed noise with zero-centered mean value, as shown in Eq. 8. It’s worth noting that, as shown in Fig. 4(b), MC can mitigate the misdirection of the incorrect box annotations and eliminate erroneous thorn-like regions.
(2) By comparing the first and third rows in Table 2, we find that using LC alone brings only limited improvement and fails to address erroneous thorn-like regions. This is because without MC in the training phase, the model is easily fitted to the train-set, having the predicted masks touch the noisy GT boxes, so the LC strategy is unable to correct the noisy GT boxes to improve their tightness.
(3) By comparing the last three rows in Table 2, we can conclude that MC and LC are not mutually excluded, and the combination of them can achieve the best performance (see the last three columns in Fig. 4). This is benefit from that MC-guided predictions can help LC to improve the tightness of boxes, which in turn can guide more accurate predictions.
4.4.2 Different Strategies for Unconfidnt Region.
@@
MonoBox essentially defines the unconfident region and improves the reliability of constraint in this region through monotonicity constraint (MC). Likewise, there are also some strategies that address the constraint of the unconfident region, such as excluding this region in loss calculation, or computing the loss using soft labels. To further verify the effectiveness of MC, we conduct an experiment to compare the performance of MC with two strategies, denoted as Exclusion and Soft Label, respectively. We choose 2D-Gaussian Label [26] as the implementation of Soft Label due to its wide usage. Note that, to avoid the influence of other components, we disable LC in this experiment. The quantitative comparison results are presented in Table 3. As can be seen, using Exclusion hardly improves the performance of the baseline. This is because simply excluding the unconfident regions could waste the possibly carried valuable information. Using Soft Label can improve the performance compared to LB, but there is still a large gap compared to UB. In contrast, our MC significantly improves the baseline’s noise tolerance, and achieves better performance than both strategies by remarkable margins.
@@
4.4.3 Unconfident Scale ($\lambda$).
The unconfident scale $\lambda$ in Eq. 3 is a heuristically determined hyperparameter related to noise level. If $\lambda$ is set higher than real noise level, the unconfident region will be overstimated and the confident region shrinks, leading to a decrease in the efficiency of correct supervision. If $\lambda$ is set lower than real noise level , the unconfident region will be understimated, leading to performance degradation due to the incorrect supervisions. We compare the variants of MonoBox using different unconfident scales $\lambda$, and the results are listed in Table 4. As can be seen, MonoBox using $\lambda=0.2$ shows the best performance on both datasets except for the metric of HD on the Synthetic noisy dataset.
Table 4: Ablation on the uncofident scale $\lambda$
$\lambda$ Real Synthetic
Dice IoU HD(px) Dice IoU HD(px)
0.1 0.836 0.742 1.985 0.795 0.706 3.327
0.2 0.849 0.764 1.920 0.803 0.714 3.338
0.3 0.844 0.760 1.964 0.794 0.705 3.468
0.4 0.828 0.733 2.011 0.782 0.688 3.534
4.5 Performance under Different Levels of Noise
@@
Figure 5: Dice of different methods under the different levels of noise in the Synthetic noisy dataset.
We conduct an experiment to analyze the stability of MonoBox and the previous anit-noise SOTAs [40, 21, 34, 44] against different level of synthetic box noises. Specifically, taking IBoxCLA as the box-supervised segmentation backbone, we train the methods multiple times using four Synthetic noisy datasets with different noise levels ($\sigma=0.1,0.2,0.3,0.4$).
@@
Fig. 5 illustrates the performance trends in terms of Dice of these methods. As can be seen, when adding a slight noise $(\sigma=0.1)$, the segmentation backbone suffers from a significant decrease in Dice from 0.827 to 0.753 (UB vs. LB). Compared to LB, NCE+RCE [21] and FSRM [44] show limited performance gains, and SSD-Det [40] and PolarT [34] even degrade the performance. In contrast, MonoBox greatly alleviates the interference of noise, improving Dice by 5.7% (0.810 vs. 0.753) compared to LB, and almost competitive to UB (0.810 vs. 0.827). As the noise level increasing, all methods show varying degrees of performance degradation. However, the curve of MonoBox is stabler and flatter compared to those of other methods, which shows that our method is more robust to noisy boxes.
5 Conclusion and Limitations
@@
In this paper, we propose MonoBox, which addresses the limitations of the existing MIL-based box-supervised segmentation method on tightness-free box annotations. Inspired by the spatial distribution of the noise, we propose monotonicity constraint (MC) to encourage a monotonicity trend that the box-inner response should be higher than the box-outer response on the unconfident regions, which provides more reliable optimizations in model training. Moreover, we propose label correction (LC), which uses the predictions guided by MC to improve the tightness of the box annotations and dynamically optimize the training process. Our approach is general and can easily cooperate with modern MIL-based box-supervised segmentation methods. The comprehensive and extensive comparisons with four anti-noise and six box-supervised segmentation state-of-the-arts (SOTA) on the public synthetic noisy dataset and the in-house real noisy dataset demonstrate that MonoBox outperforms the other SOTAs significantly, which verify the effectiveness of our method and the great application value in the realistic clinical practice.
Nevertheless, MonoBox still has some limitations, mainly in terms of empirical and static unconfident scale setting : (1) Empirical setting: Monobox requires the scale to roughly align with the noise level of the dataset, which is sometimes not that easy. (2) Static setting: we utilize a static scale for all samples during training, but each sample has different levels of noise. In the future, we will explore methods to adaptively perceive the unconfident region of each training sample.

### v5

@@ Introduction
We view MonoBox as a plug-and-play anti-noise method, and compare it with four state-of-the-art (SOTA) anti-noise methods, i.e., SSD-Det (Wu et al. 2023), NCE+RCE (Ma et al. 2020), PolarT (Wang and Xia 2022), and FSRM (Zhu et al. 2023) on the Real noisy dataset and the Synthetic noisy dataset with noise level $\sigma=0.2$. We train these methods using two box-supervised (BSS) segmentation backbones, i.e., WeakPolyp (Wei et al. 2023) and IBoxCLA (Wang et al. 2023). Note that, we denote UB and LB as backbones trained using the clean and noisy datasets, respectively, and since there are no clean annotations in the Real noisy dataset, we do not report results of UB for the Real noisy dataset.
The comparison results are provided in Table 1. Comparing the results of LB and UB of the two backbones, we can see that on Synthetic noisy dataset, the noise leads a decrease of Dice by 7.4% and 9.2% for WeakPolyp and IBoxCLA, respectively. This exhibits the high sensitivity of the existing BSS methods to noisy boxes. When using MonoBox for optimization, the performance margin compared to the UB is significantly narrowed for both backbones. Specifically, Dice of IBoxCLA is improved by 4.6% and 6.8% , and Dice of WeakPolyp is improved by 2.9% and 6.3%, on the Real and Synthetic noisy datasets, respectively.
Moreover, our method achieves the best performance among all anti-noise methods. IBoxCLA with our method exceeds that with the second-best method, i.e., FSRM, by 3.3% and 5.5% in Dice on the Real and Synthetic noisy datasets, respectively. WeakPolyp with our method outperforms that with FSRM by 1.9% and 4.1% in Dice on the Real and Synthetic noisy datasets, respectively. This is because that FSRM aims to correct the original noisy boxes using the model’s predictions, but the predictions guided by traditional MIL are unreliable under the supervision of noisy labels. In contrast, the monotonicity constraint of MonoBox can provide reliable supervision even under noisy labels.
Figure 4 shows the visualization results of different methods using IBoxCLA as the backbone. It can be seen that the model trained with our method can accurately find the inconspicuous polyps and segment more reasonable boundaries.
Ablation Study
Effectiveness of the key components.
@@
To verify the effectiveness two key components of MonoBox, i.e., monotonicity constraint (MC) and label correction (LC), we train three variants of MonoBox by disabling MC and/or LC, and IBoxCLA is used as the BSS backbone. The segmentation results are presented in Table 2 and Figure 5. Based on these results, two key conclusions can be made as follows:
@@
(1) By comparing the first two rows in Table 2 corresponding to (a) and (b) in Figure 5, we can see that MC significantly improves the tolerance of the model to noisy box annotations. From Figure 5(a), we can see that on the Real noisy dataset, the baseline tends to produce small thorn-like regions (indicated by yellow arrows) at the boundary. This is because the doctors often annotated over-sized boxes, and thus the model has to output splinters to touch the incorrect boundaries. This phenomenon does not occur in the Synthetic noisy dataset, because we use gaussian distributed noise with zero-centered mean value, as shown in Eq. 8. It’s worth noting that, as shown in Figure 5(b), MC can mitigate the misdirection of the incorrect box annotations and eliminate erroneous thorn-like regions.
(2) By comparing the first and third rows in Table 2 and seeing Figure 5(c), we find that using LC alone brings only limited improvement and fails to address erroneous thorn-like regions. However, when LC is combined with MC, it can lead more improvement and eliminate the thorn-like regions, as shown in Figure 5(d). To clearly show the mechanism of LC, we visualize the label accuracy curve with/without MC across the training process on the Synthetic noisy dataset in Figure 6. Without MC, LC hardly improves the accuracy of labels, while with MC, LC significantly and continuously improves the label accuracy. This is because MC optimizes the constraint in unconfident regions, having the predicted masks more precisely, which can guide LC to improve the tightness of boxes, which in turn can guide more accurate predictions. Therefore, MC and LC complement each other and the combination of them achieves the best performance.
Figure 6: Label accuracy curve across the training process on the Synthetic noisy dataset. The ordinate represents the IoU between the training box label and the clean box label. Following the implementation details, the label correction is performed every 10 epochs.
Different Strategies for Unconfident Region.
@@
MonoBox essentially defines the unconfident region and improves the reliability of constraint in this region through monotonicity constraint (MC). Likewise, there are also some strategies that address the constraint of the unconfident region, such as excluding this region in loss calculation, or computing the loss using soft labels. To further verify the effectiveness of MC, we conduct an experiment to compare the performance of MC with two strategies, denoted as Exclusion and Soft Label, respectively. We choose 2D-Gaussian Label (Qadir et al. 2021) as the implementation of Soft Label due to its wide usage. Note that, to avoid the influence of other components, we disable LC in this experiment. The quantitative comparison results are presented in Table 3. As can be seen, using Exclusion hardly improves the performance of the baseline. This is because simply excluding the unconfident regions could waste the possibly carried valuable information. Using Soft Label can improve the performance compared to LB, but there is still a large gap compared to UB. In contrast, our MC significantly improves the baseline’s noise tolerance, and achieves better performance than both strategies by remarkable margins.
@@
Hyperparameters Choices.
Table 3 compares our results on the Real noisy dataset with different choices of hyperparameters: confindent scale $\lambda$ in Eq. 3, IoU threshold $\tau$ and interval epoch $t$ in label correction. As can be seen, MonoBox using $\lambda=0.2$, $\tau=0.7$, and $t=10$ shows the best performance, respectively. Note that, MonoBox does not rely on hyperparameter tuning when applied to different datasets. As shown in Figure 7, we keep hyperparameter choices constant ($\lambda=0.2$, $\tau=0.7$, $t=10$) on datasets with different noise levels, MonoBox can maintain stable performance and outperform other SOTA methods.
Table 4: Ablation on $\lambda$, $\tau$ and $t$ on the Real noisy dataset
$\lambda$ Dice HD(px) $\tau$ Dice HD(px) $t$ Dice HD(px)
0.1 0.836 2.108 0.5 0.842 1.994 2 0.840 2.057
0.2 0.849 1.920 0.6 0.846 1.953 5 0.844 1.987
0.3 0.844 1.964 0.7 0.849 1.920 10 0.849 1.920
0.4 0.838 2.011 0.8 0.845 1.988 20 0.842 2.043
Performance under Different Levels of Noise
@@
Figure 7: Dice of different methods under the different levels of noise on the Synthetic noisy dataset.
@@
We conduct an experiment to analyze the stability of MonoBox and the previous anit-noise SOTAs against different level of synthetic box noises. Specifically, taking IBoxCLA as the BSS backbone, we train the methods multiple times using four Synthetic noisy datasets with different noise levels $\sigma$. Figure 7 illustrates the performance trends in terms of Dice of these methods. As can be seen, when adding a slight noise $(\sigma=0.1)$, the segmentation backbone suffers from a significant decrease in Dice from 0.827 to 0.753 (UB vs. LB). Compared to LB, NCE+RCE (Ma et al. 2020) and FSRM (Zhu et al. 2023) show limited performance gains, and SSD-Det (Wu et al. 2023) and PolarT (Wang and Xia 2022) even degrade the performance. In contrast, MonoBox greatly alleviates the interference of noise, improving Dice by 5.7% (0.810 vs. 0.753) compared to LB, and almost competitive to UB (0.810 vs. 0.827). As the noise level increasing, all methods show varying degrees of performance degradation. However, the curve of MonoBox is stabler and flatter compared to those of other methods, which shows that our method is more robust to noisy boxes.
Generality for other scenario and method
MonoBox is not only applicable to any scenario, but also can be easily implemented with other MIL-based BSS methods. To verify this, we conduct an experiment on COCO (Lin et al. 2014), set BoxInst (Tian et al. 2021) as the backbone, and expand the MonoBox to adapt its MIL-based loss, i.e., projection loss. The tightness-free box annotations are generated by Eq.  8. As shown the results in Table 5, MonoBox significantly narrows the performance gap between LB and UB and outperforms the second best anti-noise method, i.e, FSRM, by a large margin, which verifies that generality of MonoBox on natural scene and other MIL-based box-supervised segmentation methods.
Table 5: Comparison results on the COCO test-dev split.
Backbone Methods AP AP50 AP75 APS APM APL
BoxInst UB 0.321 0.551 0.324 0.156 0.343 0.435
LB 0.248 0.480 0.225 0.117 0.282 0.329
FSRM 0.259 0.488 0.240 0.124 0.293 0.344
MonoBox(Ours) 0.308 0.532 0.304 0.149 0.331 0.417
Conclusion and Future Work
@@
In this paper, we propose MonoBox, which addresses the limitations of the existing BSS method on tightness-free box annotations. Inspired by the specific spatial distribution of the noise, we propose MC to provide more reliable optimizations in model training. Moreover, we propose LC to improve the tightness of the box annotations and dynamically optimize the training process. Our approach is general and can easily cooperate with modern MIL-based BSS methods. The comprehensive experiments on the public synthetic noisy dataset and the in-house real noisy dataset demonstrate that MonoBox can effectively solve the realistic problems in clinical practice and has great application value. Nevertheless, MonoBox currently adopts a uniform unconfident scale for all samples during training, we will explore methods to adaptively perceive the appropriate unconfident scale of each training sample in the future.
Acknowledgments
This work was supported in part by National Key R&D Program of China (Grant No. 2023YFC2414900), Key R&D Program of Hubei Province of China (No.2023BCB003), Fundamental Research Funds for the Central Universities (2021XX-JS033), Wuhan United Imaging Healthcare Surgical Technology Co., Ltd.